### v1

@@ 5.3 Soil Contamination Analysis
Similarly, [Egypt_contamnation_2022] employed ordinary Kriging (OK) for the generation of spatial distribution maps for six heavy metals (Cr, Co, Cu, Cd, Pb, Zn). Principal component analysis (PCA) and contamination factor (CF) calculations were utilized to assess soil contamination severity in the El-Moheet drainage basin, located on the western bank of the Nile River in El-Minia Governorate, Egypt.
@@
An ongoing study of Lat et al.[Malysia_Contamination_soil_2024] performed a detailed geochemical assessment of soil contamination within the Pasir Gudang industrial region of Johor, Malaysia, aiming to characterize the spatial distribution and contamination intensity of selected heavy metals across diverse land-use categories. Topsoil samples (0–20 cm depth) were systematically collected from twenty sites representing industrial, residential, riparian, and educational zones. The samples were subjected to quantitative analysis for major heavy metals, specifically copper (Cu), chromium (Cr), nickel (Ni), zinc (Zn), and lead (Pb), utilizing inductively coupled plasma mass spectrometry (ICP-MS). Contamination levels were quantitatively evaluated using the geoaccumulation index ($I_{g}eo$) and contamination factor ($C_{f}$) metrics. The results indicated that most heavy metal concentrations were below established guideline limits, suggesting the area is generally uncontaminated or only slightly contaminated. The values of $I_{g}eo$ and $C_{f}$ supported these findings, indicating low levels of contamination and minimal anthropogenic impact on soil quality. However, marginally higher concentrations were detected near certain industrial operations, reflecting localized pollution. The study concluded that current soil management practices in Pasir Gudang current practices were effective, but emphasized the need for ongoing monitoring is essential to detect and mitigate potential pollutant accumulation in industrialized sectors. It is imperative to note that this research methodology does not encompass the application or integration of sophisticated statistical modeling or machine learning (ML) algorithms in its analysis.
@@
A recent study by [Portugal_contaminationAgroSoil_2025] introduced an optimized methodology for the spatial assessment of heavy metal contamination risk, circumventing the logistical constraints associated with physical soil sampling. The study formulated a Probabilistic Pollution Index (PPI) through the integration of Geographic Information System (GIS) technologies with an eight-parameter probability-risk matrix. The parameters incorporated into the model included proximity to roads, industrial sites, soil pH, soil organic matter content, terrain slope, soil texture, proximity to mining areas, and drainage potential. Each parameter underwent a classification and reclassification process to generate a contamination risk surface, whereby individual pixels were categorized into five distinct risk tiers. The efficacy of the PPI was evaluated through comprehensive data analysis involving parameter classification, reclassification procedures, and risk mapping validation.
@@ 5.4 Soil Salinization
Similarly, [li_spatiotemporal_2024] devised a methodology for salinization monitoring in arid regions utilizing multispectral Landsat imagery spanning $2001$ to $2021$ in the Yutian Oasis, China. The study quantitatively assessed the classification accuracy of Support Vector Machine (SVM) and Classification and Regression Tree (CART) algorithms for high-resolution spatiotemporal delineation of soil salinity distribution, providing robust analytical tools to inform salinization mitigation strategies and promote sustainable agricultural land use planning.
@@ 5.5 Land Use Land Cover Change
Land use and land cover (LULC) change constitutes a critical domain within spatial analysis, emphasizing the quantification and characterization of landscape modifications resulting from anthropogenic activities and natural dynamics. This process is influenced by a range of drivers, including urban expansion, agricultural intensification, and socio-economic development.
Numerous research efforts have employed remote sensing imagery and geospatial analytical methods to systematically detect, monitor, and evaluate LULC transitions across various geographic regions. In [shakya_land_2023], land use and cover changes in the Guntur district of Singapore were analyzed from 2014 to 2021 using Principal Component Analysis (PCA), K-means clustering, and the MAD algorithm. Their findings highlighted the impacts of water scarcity on cropping patterns, land use dynamics, rainfall variability, biodiversity risks, and industrial development, with notable changes observed in horticultural land, built-up areas, harvested land, and wasteland.
@@
In [brito_land_2025], analyzed land use and land cover (LULC) changes in the the Portal do Sertão situated in Bahia State, Brazil, spanning the temporal interval from 1985 to 2022. The primary objective was to quantify and characterize the nature of spatial transformations and their underlying factors. The methodological framework integrated advanced remote sensing (RS) data acquisition using MapBiomas Collection 8.0 dataset, and geographic information system (GIS) to reclassify the land‐cover into four categories: Forest Formation, Agriculture, Urban Area and Water Bodies. These classifications were delineated across discrete temporal snapshots corresponding to the years 1985, 1990, 1995, 2000, 2005, 2010, 2015, 2020, 2022. Their results revealed significant reductions in agricultural and livestock-associated land extents, concomitant with dramatic increase of urbanized zones and expansion of water surfaces. The spatiotemporal variability of forest cover manifested as episodic fluctuations and were negatively impacted by urbanization and agricultural pressure. The authors inferred that the integration of remote sensing datasets with GIS analytical tools proves efficacious in tracking and modeling spatial-temporal LULC modifications, thereby furnishing critical spatial datasets to inform territorial planning initiatives and environmental policymaking within the region. The investigation utilized categorical spatial analytical methodologies, specifically employing transition matrix formulations and quantification of areal modifications, supplemented by descriptive statistical summaries. No implementation of inferential statistical hypothesis testing, spatial interpolation techniques, or machine learning algorithms was conducted within the scope of the study.
@@
Furthermore, [kucuk_analysis_2025] conducted a comprehensive spatio-temporal analysis of land cover and land use changes across seven geographic regions of the Republic of Türkiye, using the CORINE (Coordination of Information on the Environment) land cover classification datasets corresponding to the temporal points of 1990, 2000, 2006, 2012, and 2018. The primary objective was regional differences in land transformation patterns quantify the extent of change within principal land cover classes including artificial surface, agricultural land uses, forested and semi-natural vegetation, wetlands, and surface water bodies. The pipeline processed and analyzed the CORINE data using Geographic Information Systems (GIS) to generate regional land-cover maps and compute the areal extent and percentage change of each category over time using quantitative metrics, such as area extent (in km²) and percentage change relative to baseline years. The analytical outcomes indicated statistically significant expansions in artificial and urbanized land cover categories, concomitant with a discernible attrition of forested and semi-natural ecosystems across the majority of the study area. Agricultural lands exhibited variable trends, with notable conversions to built-up zones and industrial areas. The study relied on GIS-based spatial quantification and descriptive analysis and did not employ statistical inference, interpolation, or machine learning methods.
@@
In the study conducted by [senadi_spatio-temporal_2025], the researchers aimed to delineate and quantify the spatiotemporal changes from agricultural to urban landscapes in the Mitidja Plain of Algeria. They employed multitemporal spectral data derived from Landsat and Sentinel-2 satellite datasets, covering the temporal interval from the year 1983 through 2023, which were processed and analyzed via the Google Earth Engine geospatial analysis platform. Two classification algorithms are investigated, including Random Forest (RF) and Smile CART (Classification and Regression Trees). An evaluative comparative analysis indicated that the RF classifier demonstrated superior performance metrics, including enhanced classification accuracy, in detecting and delineating land cover transition zones marked by urban expansion subsequent to agricultural land use. The quantitative results of their analyses revealed a persistent reduction in the extent of cultivated agricultural lands corresponding to an approximate total loss of 28.04% of initial cultivated land area. Conversely, urbanized, built-up land cover exhibited an increasing by 25.67% relative to initial urban extents during the same time frame.
@@ 5.6 Land Degradation Assessment
Land degradation [Barman2020] refers to the decline in land quality and productivity, primarily due to natural processes, like erosion, salinization, and drought; human activities such as deforestation, overgrazing, and poor irrigation practices. It affects soil health, vegetation cover, water resources, and overall land functionality. Its categorized on four types : physical degradation(soil compaction, erosion, crusting), chemical degradation (salinity, sodicity, nutrient depletion), biological degradation (loss of vegetation cover, biodiversity loss) and anthropogenic degradation (unsustainable agriculture, urban sprawl, deforestation).
@@
An alternate investigation, [Blanche2024] evaluated land cover degradation caused by semi-mechanised and artisanal mining activities in the Mbale, region of Cameroon. This assessment used an integrated remote sensing and digital photogrammetric approach, utilizing multispectral imagery acquired from Sentinel-2 satellite sensors for the years 2019, 2021, and 2023. The collected satellite data underwent pre-processing workflows that included atmospheric correction and seasonal normalization through the computation of spectral vegetation and moisture indices including the Normalized Difference Vegetation Index (NDVI), Normalized Difference Water Index (NDWI), Brightness Index, and Soil Crust Index, to mitigate confounding effects of atmospheric variability and phenological cycles. Subsequently, supervised classification algorithms, namely, Maximum Likelihood Classification (MLC) were implemented to categorize land cover classes, which encompassed vegetative cover, bare soil substrata, hydrological features, and anthropogenic disturbed zones indicative of mining activity. This research underscores the efficacy of combining remote sensing-derived spectral indices, advanced photogrammetric techniques, and field validation protocols to monitor the environmental repercussions of mining activities. Moreover, it contributes to the empirical foundation necessary for informing sustainable land management strategies in environmentally sensitive, mining-impacted landscapes.
@@ 5.7 Climate Change
Table 3: Comparative analysis of application papers in geospatial soil quality analysis systems
@@
El Behairy et al. (2024) [el_behairy_accurate_2024] Predict Soil Quality Index (SQI) ANN (no GIS) using 16 soil attributes Local (Drylands) Pure ML with lab-based data, no spatial mapping
@@ 6 Challenges and Limitations
Despite advancements in soil quality analysis, several challenges and limitations remain. Even though geospatial techniques, machine learning, remote sensing, and digital soil mapping have greatly advanced, they still face substantial challenges. The following highlights some of the key issues documented in recent literature.
Many regions, especially in developing countries, suffer from inadequate field data, which poses significant challenges for training and validating models [folorunso_exploring_2023, DeCaires2025]. Ground-truth data in these areas are often sparse, outdated, or unevenly distributed, limiting the calibration and validation of models [AppChallgsDSMAfrica2024]. Additionally, biological soil properties such as microbial biomass, enzyme activity, and soil fauna are often underrepresented in Digital Soil Mapping (DSM) frameworks due to the complexity and cost of measurement [DevSQAframework2025]. The problem is intensified by variations in the soil classification systems, measurement protocols, units and scales used in geographic regions and research studies, which complicate the integration and comparability of data [Lagacherie2025]. There is also a semantic gap in how terms like "soil quality" and "soil health" are conceptualized across disciplines [DevSQAframework2025]. Furthermore, many models do not report or propagate prediction uncertainty, making users unaware of the confidence levels in spatial outputs [DeCaires2025]. The management, storage, processing, and visualization of large, multi-temporal, multi-source geospatial datasets can be computationally intensive, especially in data-poor or low-resource environments, where challenges such as reproducible workflows, automation, and insufficient data infrastructure further complicate the process [DeCaires2025, Liakos2022]. Machine learning models, which are frequently perceived as "black boxes," also face challenges related to limited interpretability, impacting decision-makers’ capacity to comprehensively understand the results [DeCaires2025]. In addition, soil quality is dynamic, with changes driven by land management, erosion, and climate variability, but many soil quality maps are static snapshots. Regular updates to these maps can be costly or restricted by a lack of up-to-date covariates and field data [AppChallgsDSMAfrica2024]. Finally, soil data and tools are often not translated into user-friendly formats for local farmers or policymakers, reducing their accessibility and practical utility.
@@ 7 Trends and Future Directions
Furthermore, the development of Explainable AI (XAI) techniques can enhance model transparency, thereby making decision-support systems more interpretable and accessible to stakeholders [Minasny2024, DeCaires2025]. Participatory sensing and citizen science initiatives—such as mobile applications, community-based monitoring, and the use of low-cost sensors—provide valuable opportunities to expand data collection efforts. Additionally, cloud computing and big data integration are crucial for processing large-scale datasets in near real time, particularly in remote or resource-limited regions. At the policy level, designing systems that directly support sustainable land management and agricultural extension programs is essential to enable more informed and effective decision-making [DeCaires2025].
@@ 8 Conclusion
Future development efforts may focus on dynamic spatio-temporal modeling, the integration of biological indicators, progress in explainable artificial intelligence (XAI) and uncertainty quantification, and the integration of Internet of Things (IoT) sensor networks with cloud-based Geographic Information System (GIS) platforms. Furthermore, adopting participatory approaches and developing policy-oriented tools are essential to ensure geospatial soil quality systems are user-centric and directly relevant to sustainable land management practices.
@@
The authors received no funding for this work.

### v2

@@ 5.3 Soil Contamination Analysis
Similarly, [Egypt_contamnation_2022] employed ordinary Kriging (OK) for the generation of spatial distribution maps for six heavy metals (Cr, Co, Cu, Cd, Pb, Zn). PCA and contamination factor (CF) calculations were utilized to assess soil contamination severity in the El-Moheet drainage basin, located on the western bank of the Nile River in El-Minia Governorate, Egypt.
@@
An ongoing study of Lat et al.[Malysia_Contamination_soil_2024] performed a detailed geochemical assessment of soil contamination within the Pasir Gudang industrial region of Johor, Malaysia, aiming to characterize the spatial distribution and contamination intensity of selected heavy metals across diverse land-use categories. Topsoil samples (0–20 cm depth) were systematically collected from twenty sites representing industrial, residential, riparian, and educational zones. The samples were subjected to quantitative analysis for major heavy metals, specifically copper (Cu), chromium (Cr), nickel (Ni), zinc (Zn), and lead (Pb), utilizing inductively coupled plasma mass spectrometry (ICP-MS). Contamination levels were quantitatively evaluated using the geoaccumulation index ($I_{geo}$) and contamination factor ($C_{f}$) metrics. The results indicated that most heavy metal concentrations were below established guideline limits, suggesting the area is generally uncontaminated or only slightly contaminated. The values of $I_{geo}$ and $C_{f}$ supported these findings, indicating low levels of contamination and minimal anthropogenic impact on soil quality. However, marginally higher concentrations were detected near certain industrial operations, reflecting localized pollution. The study concluded that current soil management practices in Pasir Gudang were effective, but emphasized the need for ongoing monitoring is essential to detect and mitigate potential pollutant accumulation in industrialized sectors. It is imperative to note that this research methodology does not encompass the application or integration of sophisticated statistical modeling or ML algorithms in its analysis.
@@
A recent study by [Portugal_contaminationAgroSoil_2025] introduced an optimized methodology for the spatial assessment of heavy metal contamination risk, circumventing the logistical constraints associated with physical soil sampling. The study formulated a Probabilistic Pollution Index (PPI) through the integration of GIS technologies with an eight-parameter probability-risk matrix. The parameters incorporated into the model included proximity to roads, industrial sites, soil pH, soil organic matter content, terrain slope, soil texture, proximity to mining areas, and drainage potential. Each parameter underwent a classification and reclassification process to generate a contamination risk surface, whereby individual pixels were categorized into five distinct risk tiers. The efficacy of the PPI was evaluated through comprehensive data analysis involving parameter classification, reclassification procedures, and risk mapping validation.
@@ 5.4 Soil Salinization
Similarly, [li_spatiotemporal_2024] devised a methodology for salinization monitoring in arid regions utilizing multispectral Landsat imagery spanning $2001$ to $2021$ in the Yutian Oasis, China. The study quantitatively assessed the classification accuracy of SVM and Classification and Regression Tree (CART) algorithms for high-resolution spatiotemporal delineation of soil salinity distribution, providing robust analytical tools to inform salinization mitigation strategies and promote sustainable agricultural land use planning.
@@ 5.5 Land Use Land Cover Change
LULC change constitutes a critical domain within spatial analysis, emphasizing the quantification and characterization of landscape modifications resulting from anthropogenic activities and natural dynamics. This process is influenced by a range of drivers, including urban expansion, agricultural intensification, and socio-economic development.
Numerous research efforts have employed remote sensing imagery and geospatial analytical methods to systematically detect, monitor, and evaluate LULC transitions across various geographic regions. In [shakya_land_2023], land use and cover changes in the Guntur district of Singapore were analyzed from 2014 to 2021 using PCA, K-means clustering, and the Multivariate Alteration Detection (MAD) algorithm. Their findings highlighted the impacts of water scarcity on cropping patterns, land use dynamics, rainfall variability, biodiversity risks, and industrial development, with notable changes observed in horticultural land, built-up areas, harvested land, and wasteland.
@@
In [brito_land_2025], researchers analyzed LULC changes in the Portal do Sertão situated in Bahia State, Brazil, spanning the temporal interval from 1985 to 2022. The primary objective was to quantify and characterize the nature of spatial transformations and their underlying factors. The methodological framework integrated advanced remote sensing (RS) data acquisition using MapBiomas Collection 8.0 dataset, and GIS to reclassify the land‐cover into four categories: Forest Formation, Agriculture, Urban Area and Water Bodies. These classifications were delineated across discrete temporal snapshots corresponding to the years 1985, 1990, 1995, 2000, 2005, 2010, 2015, 2020, 2022. Their results revealed significant reductions in agricultural and livestock-associated land extents, concomitant with dramatic increase of urbanized zones and expansion of water surfaces. The spatiotemporal variability of forest cover manifested as episodic fluctuations and were negatively impacted by urbanization and agricultural pressure. The authors inferred that the integration of remote sensing datasets with GIS analytical tools proves efficacious in tracking and modeling spatial-temporal LULC modifications, thereby furnishing critical spatial datasets to inform territorial planning initiatives and environmental policymaking within the region. The investigation utilized categorical spatial analytical methodologies, specifically employing transition matrix formulations and quantification of areal modifications, supplemented by descriptive statistical summaries. No implementation of inferential statistical hypothesis testing, spatial interpolation techniques, or machine learning algorithms was conducted within the scope of the study.
@@
Furthermore, [kucuk_analysis_2025] conducted a comprehensive spatio-temporal analysis of land cover and land use changes across seven geographic regions of the Republic of Türkiye, using the CORINE (Coordination of Information on the Environment) land cover classification datasets corresponding to the temporal points of 1990, 2000, 2006, 2012, and 2018. The primary objective was regional differences in land transformation patterns quantify the extent of change within principal land cover classes including artificial surface, agricultural land uses, forested and semi-natural vegetation, wetlands, and surface water bodies. The pipeline processed and analyzed the CORINE data using GIS to generate regional land-cover maps and compute the areal extent and percentage change of each category over time using quantitative metrics, such as area extent (in km²) and percentage change relative to baseline years. The analytical outcomes indicated statistically significant expansions in artificial and urbanized land cover categories, concomitant with a discernible attrition of forested and semi-natural ecosystems across the majority of the study area. Agricultural lands exhibited variable trends, with notable conversions to built-up zones and industrial areas. The study relied on GIS-based spatial quantification and descriptive analysis and did not employ statistical inference, interpolation, or machine learning methods.
@@
In the study conducted by [senadi_spatio-temporal_2025], the researchers aimed to delineate and quantify the spatiotemporal changes from agricultural to urban landscapes in the Mitidja Plain of Algeria. They employed multitemporal spectral data derived from Landsat and Sentinel-2 satellite datasets, covering the temporal interval from the year 1983 through 2023, which were processed and analyzed via the Google Earth Engine geospatial analysis platform. Two classification algorithms are investigated, including RF and Smile CART (Classification and Regression Trees). An evaluative comparative analysis indicated that the RF classifier demonstrated superior performance metrics, including enhanced classification accuracy, in detecting and delineating land cover transition zones marked by urban expansion subsequent to agricultural land use. The quantitative results of their analyses revealed a persistent reduction in the extent of cultivated agricultural lands corresponding to an approximate total loss of 28.04% of initial cultivated land area. Conversely, urbanized, built-up land cover exhibited an increasing by 25.67% relative to initial urban extents during the same time frame.
@@ 5.6 Land Degradation Assessment
Land degradation [Barman2020] refers to the decline in land quality and productivity, primarily due to natural processes, like erosion, salinization, and drought; human activities such as deforestation, overgrazing, and poor irrigation practices. It affects soil health, vegetation cover, water resources, and overall land functionality. It is categorized in four types : physical degradation(soil compaction, erosion, crusting), chemical degradation (salinity, sodicity, nutrient depletion), biological degradation (loss of vegetation cover, biodiversity loss) and anthropogenic degradation (unsustainable agriculture, urban sprawl, deforestation).
@@
An alternate investigation, [Blanche2024] evaluated land cover degradation caused by semi-mechanised and artisanal mining activities in the Mbale, region of Cameroon. This assessment used an integrated remote sensing and digital photogrammetric approach, utilizing multispectral imagery acquired from Sentinel-2 satellite sensors for the years 2019, 2021, and 2023. The collected satellite data underwent pre-processing workflows that included atmospheric correction and seasonal normalization through the computation of spectral vegetation and moisture indices including the NDVI, Normalized Difference Water Index (NDWI), Brightness Index, and Soil Crust Index, to mitigate confounding effects of atmospheric variability and phenological cycles. Subsequently, supervised classification algorithms, namely, Maximum Likelihood Classification (MLC) were implemented to categorize land cover classes, which encompassed vegetative cover, bare soil substrata, hydrological features, and anthropogenic disturbed zones indicative of mining activity. This research underscores the efficacy of combining remote sensing-derived spectral indices, advanced photogrammetric techniques, and field validation protocols to monitor the environmental repercussions of mining activities. Moreover, it contributes to the empirical foundation necessary for informing sustainable land management strategies in environmentally sensitive, mining-impacted landscapes.
@@ 5.7 Climate Change
Table 3: Recapitulation of application papers in geospatial soil quality analysis systems
@@
El Behairy et al. (2024) [el_behairy_accurate_2024] Predict SQI ANN (no GIS) using 16 soil attributes Local (Drylands) Pure ML with lab-based data, no spatial mapping
@@ 6 Challenges and Limitations
Despite advancements in soil quality analysis, several challenges and limitations remain. Even though geospatial techniques, machine learning, remote sensing, and digital soil mapping have greatly advanced, they still face substantial challenges. The following highlight some of the key issues documented in recent literature.
Many regions, especially in developing countries, suffer from inadequate field data, which poses significant challenges for training and validating models [folorunso_exploring_2023, DeCaires2025]. Ground-truth data in these areas are often sparse, outdated, or unevenly distributed, limiting the calibration and validation of models [AppChallgsDSMAfrica2024]. Additionally, biological soil properties such as microbial biomass, enzyme activity, and soil fauna are often underrepresented in DSM frameworks due to the complexity and cost of measurement [DevSQAframework2025]. The problem is intensified by variations in the soil classification systems, measurement protocols, units and scales used in geographic regions and research studies, which complicate the integration and comparability of data [Lagacherie2025]. There is also a semantic gap in how terms like "soil quality" and "soil health" are conceptualized across disciplines [DevSQAframework2025]. Furthermore, many models do not report or propagate prediction uncertainty, making users unaware of the confidence levels in spatial outputs [DeCaires2025]. The management, storage, processing, and visualization of large, multi-temporal, multi-source geospatial datasets can be computationally intensive, especially in data-poor or low-resource environments, where challenges such as reproducible workflows, automation, and insufficient data infrastructure further complicate the process [DeCaires2025, Liakos2022]. Machine learning models, which are frequently perceived as "black boxes," also face challenges related to limited interpretability, impacting decision-makers’ capacity to comprehensively understand the results [DeCaires2025]. In addition, soil quality is dynamic, with changes driven by land management, erosion, and climate variability, but many soil quality maps are static snapshots. Regular updates to these maps can be costly or restricted by a lack of up-to-date covariates and field data [AppChallgsDSMAfrica2024]. Finally, soil data and tools are often not translated into user-friendly formats for local farmers or policymakers, reducing their accessibility and practical utility.
@@ 7 Trends and Future Directions
Furthermore, the development of XAI techniques can enhance model transparency, thereby making decision-support systems more interpretable and accessible to stakeholders [Minasny2024, DeCaires2025]. Participatory sensing and citizen science initiatives—such as mobile applications, community-based monitoring, and the use of low-cost sensors—provide valuable opportunities to expand data collection efforts. Additionally, cloud computing and big data integration are crucial for processing large-scale datasets in near real time, particularly in remote or resource-limited regions. At the policy level, designing systems that directly support sustainable land management and agricultural extension programs is essential to enable more informed and effective decision-making [DeCaires2025].
@@ 8 Conclusion
Future development efforts may focus on dynamic spatio-temporal modeling, the integration of biological indicators, progress in XAI and uncertainty quantification, and the integration of Internet of Things (IoT) sensor networks with cloud-based GIS platforms. Furthermore, adopting participatory approaches and developing policy-oriented tools are essential to ensure geospatial soil quality systems are user-centric and directly relevant to sustainable land management practices.
@@
The authors declare that they received no funding for this work.